OpenStudy (theloshua):

25x^2 - 40x +16 = 0 solve

Vocaloid (vocaloid):

that factors to (5x-4)^2 = 0

Vocaloid (vocaloid):

now solve for x

OpenStudy (anonymous):

Follow order of operations. 25x^2 = 625. 625x - 40x + 16 = 0. Next, subtract the same varibles. 625x - 40x =585x +16 = 0. Then, make both sides equal. 585x +16 -16 =0 -16. 585x = -16. Finally, 585x/585 = -16/585. x = -0.27.

Vocaloid (vocaloid):

^uh... not quite...

Vocaloid (vocaloid):

to solve (5x-4)^2 = 0, we take the square root of both sides 5x - 4 = 0 then add 4 to both sides 5x = 4 x = 4/5